Question


(1.)


130Cm of a gassy 20 degree celsiusi exert a pressure of 750mmHg. Calculate It's Pressure if it's Volume is increased to 150cm at 35 degree celsiusi ?



Question


(2.)


To what temperature must a given mass of Nitrogen at 0 degree celsiusi be heated so that both it's Volume and Pressure will be doubled ?

1
Expert's answer
2022-02-02T22:01:03-0500

Question 1


Unit conversions:

130 cm3 = 130 mL = 0.130 L

20oC = 273 + 20 = 293 K

150 cm3 = 150 mL = 0.150 L

35oC = 273 + 35 = 308 K

According to the combined gas law,

"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2};"


"P_2=\frac{P_1V_1T_2}{T_1V_2}=\frac{750\ mmHg\times0.130\ L\times308\ K}{293\ K\times0.150\ L}=683\ mmHg"


Answer: 683 mmHg


Question 2


0oC = 273 K

Let P1 and V1 be 1 atm and 1 L, respectively. Then, P2 and V2 are equal to 2 atm and 2 L, respectively. According to the combined gas law,

"\frac{1\ atm\times1\ L}{273\ K}=\frac{2\ atm\times2\ L}{x}."


Solving for x,


"x=\frac{273\ K\times2\ atm\times2L}{1\ atm\times1\ L}=1092\ K=1092-273=819\degree{C}"


Answer: 819oC